OpenStudy (anonymous):

Determine values of k, l,m and n such that the following function g(x) is continuous and differentiable at all points. (2x^2 − n if x<−2 ) g(x)=(mx + l if −2<=x<2 ) (kx^2 + 1 if x>=2 )

OpenStudy (anonymous):

any fixed ponts? like 2 or -2?

OpenStudy (anonymous):

no i put it up exactly as it is on the problem sheet im doing

OpenStudy (anonymous):

4k= m , m= -8 so k=-2 . l=-9

OpenStudy (anonymous):

n=1

OpenStudy (anonymous):

how@manorajan.

OpenStudy (anonymous):

how did you get those answers?

OpenStudy (anonymous):

differentiate & set the derivatives equal of the secon two equations. this gives 4k=m at x=2. do this to the firat two at x=-2. now set the values of the functions at 2 & -2 same. this gives k,l.m.n

OpenStudy (anonymous):

i.e 4k=m & from m=4x at x=-2 m=-8. so k must be -2. at x=-2 value of the last fn is -7. to be continuous the value of the second function must be the same. so -8x + l = -7

OpenStudy (anonymous):

differentiate 2x^2-n & set the derivative equal of the secon two equations. or differentiate 2x^2-n,and mx+l,and kx^2+1 & set the derivative equal of the secon two equations.?

OpenStudy (anonymous):

4k+1=2m+l . you know k & m from above. this gives -7=-16+l. l must be -9, no?

OpenStudy (anonymous):

you need to do it in pairs. the second two first, the first two next. then set values equal .. in pairs. at points 2 & -2. this gives you four equations & you have four unknowns. you can do it in any way of pairing as you like..

OpenStudy (anonymous):

this is still new to me.. can you give me the first two derived equations and what you put them equal to plz?

OpenStudy (anonymous):

ok i got it now,,,,the only thing im not sure of is when to use x=-2 and x=2?

OpenStudy (anonymous):

right. i'll write it down fully, but remember that this is just a suggestion. Most places you need to jump out & find an easier way.. differentiate the last two equations and set them equal at x=2 :: 2.k.x=m; i.e 4k=m Now, do this to the first pair, set them equal at x=-2:: 4x=m; i.e m=-8. this gives you m. use this in 4k=m .. so k=-2. now since the curve is continuous, the values of the functions at the endpoints are the same as the next fn after them. for the first two at x-2:: 8-n=-2.m+l. m=-8, so 8-n=16+l. we dont know l yet. from the second two at x=2::4.k+1=2.m+l. we know k&m. this gives us l=-9. use this in 8-n=16+l. this gives 8-n=7. so n=1

OpenStudy (anonymous):

ok i have it now thanks :)

OpenStudy (anonymous):

like i said, if your cuve is continuous then the pieces must lie end to end. so from one function to the next.. at the end points (here these were 2 & -2) their values & their derivatives must be the same. this will give you sufficient equations for all your unknowns. all you need to do is solve them.

OpenStudy (anonymous):

thanks